# MemAscend: System Memory Optimization for SSD-Offloaded LLM Fine-Tuning

### v2

@@ V Memory Efficiency and Performance Analysis
By streamlining memory consumption, the MemAscend framework reduces system memory demand relative to earlier SSD offloading techniques. Those lower system memory footprints can translate directly into cheaper training setups; in other words, the same hardware budget can now accommodate models with more trainable parameters. The lower system memory usage also unlocks longer context windows, because the larger activation-checkpoint tensors required for extended sequences can be moved from GPUs to system memory without exceeding capacity. Therefore, instead of slowing down the training process, the overall performance benefits follow as well. Meanwhile, the extra system memory headroom permits larger batch sizes, driving GPUs toward higher utilization and boosting overall throughput. The upcoming analysis examines these gains in detail and explains how the reclaimed RAM supports longer context lengths and bigger batches, both of which contribute to faster training at lower cost. These experiments are conducted using Configuration 1, as described in Table II.
@@ V-A Memory Efficiency
Figure 10 first breaks down system memory usage while training the Qwen2.5-7B model [34], highlighting substantial savings in every major component. Notably, both the compared Zero-Infinity and the proposed MemAscend are integrated with the Liger-Kernel, Flash-Attention, and offloaded gradient checkpointing for reducing residual memory usage. In the proposed MemAscend, the adaptive buffer pool is first introduced to reduce memory fragmentation. For the Qwen2.5-7B model in fp16, this approach requires about 2.46 GiB of memory for buffer allocations. In contrast, the standard SSD offloading method reserves a fixed 9.14 GiB. This change cuts memory use in the parameter buffer pool by 73%. To eliminate peak memory spikes from overflow checking, MemAscend implements the fused overflow check mechanism. Conventional checks rely on intermediate buffers, adding a 1.25$\times$ overhead and consuming 35.46 GiB. The fused mechanism maintains numerical stability without this overhead, reducing usage to 0 GiB. SSD offloading also depends on pinned memory to move data between CPU and GPU. Traditional allocation uses power-of-two alignment, introducing significant waste. MemAscend applies the alignment-free pinned memory allocation that allocates only the required memory. This reduces pinned memory overhead from 24.90 GiB to 1.63 GiB, a 93.5% decrease. Some components remain constant across methods, including the gradient flat buffer and small system allocations like optimizer state buffers and the swap-out buffer. These account for stable memory usage of roughly 28.37 GiB and 11.17 GiB, respectively. In summary, MemAscend brings total peak system memory usage down from 109.04 GiB under ZeRO-Infinity to 43.64 GiB, a 60% reduction. This allows for training larger models and handling longer input sequences within the same memory budget.
@@ VI-A Setup
All experiments were carried out on two servers. Detailed hardware specifications are provided in Table II. MemAscend is evaluated using mixed-precision training of large language models with SSD offloading. The workloads vary by model architecture, context length, and batch size. Models include Llama3.1-8B and Qwen2.5, with parameter sizes of 7 billion, 14 billion and 32 billion. All models use the Hugging Face Transformers library (v4.46.3). The context lengths range from 4,096 to 131,072 tokens. Batch sizes vary across different experiments. The main evaluation metric is peak system memory usage during training and training throughput in tokens per second to assess the overall performance.
TABLE II: Hardware Specifications for Two Configurations
Component Configuration 1 Configuration 2
CPU Intel(R) Xeon(R) 6780E 2 x AMD EPYC 7282
CPU Memory 1TB 6400 MT/s DDR5 1TB 3200 MT/s DDR4
PCIe Gen5 Gen4
GPU 2 x NVIDIA H100 PCIe w/o NVL 1 x NVIDIA A5000
SSD 1 x DapuStor Haishen5 H5100 7.5 TB 2 x Phison AI100E 2TB
To evaluate the effectiveness of MemAscend, the SOTA ZeRO-Infinity is compared, and both systems share a common setup to ensure fairness. Static memory components, including model parameters and optimizer states, are stored entirely on the SSD. To reduce GPU memory usage, gradient checkpointing, Flash-Attention, and the Liger-Kernel are enabled for both setups. For experiments that require offloaded gradient checkpointing, a modified version of Unsloth is utilized, adapted from its original single-GPU design.
VI-B Experimental Results
@@
Figure 13: Comparison of memory overhead for overflow checks between ZeRO-Infinity and MemAscend across different CPU types and various models.
@@ VI-B1 Assessment of Individual Methods
This evaluation examines the performance of the direct NVMe engine, described in Section IV-E, for SSD I/O in training scenarios with SSD offloading. The engine manages offloading of fp16 and fp32 model weights, gradients, and optimizer states. To determine suitable tensor sizes for benchmarking, model weight distributions were analyzed, excluding tensors that are too small to benefit from offloading. For instance, tensors with fewer than two million elements perform better in CPU memory. Common dense LLM architectures include several layers suitable for SSD offloading, such as feedforward layers, q, k, v, and o projection layers, the embedding layer, and the linear head layer. To ensure a precise comparison with the original SSD offloading technique, tensor size ranges for fp16 and fp32 data were collected across various models. The analysis also considers multi-GPU setups, where data are split across processes. For example, a 100-million-element tensor divides into 50 million elements per process in a two-GPU configuration. Benchmarks used Configuration 2, as specified in Table II. The original SSD offloading system, reliant on filesystem I/O, utilized two SSDs in a RAID0 array formatted with ext4. By contrast, MemAscend employs the eirect NVMe engine, which bypasses software RAID and sends requests directly to SSDs through the NVMe driver.
Figures 14(a) and 14(b) display write latency and bandwidth across tensor sizes. The direct NVMe engine consistently lowers write latency. For a 2,097,152-byte tensor, latency decreases from 988.222 microseconds to 219.418 microseconds, and for a 3,114,270,720-byte tensor, it drops from 304,604.779 to 266,212.637 microseconds. This improvement results from eliminating filesystem and RAID layer overhead. The engine also delivers higher write bandwidth across all tested sizes, achieving up to 4.5x gains and an average improvement of 72.04%. As tensor sizes increase, the performance gap narrows, likely because filesystem and RAID overhead become less significant when host computation dominates performance. Figures 14(c) and 14(d) show read latency and bandwidth. Read latency remains similar to the filesystem approach, as reads involve less overhead without costly allocation. However, the direct NVMe engine provides more consistent read bandwidth, while the filesystem-based method exhibits greater variability. The direct NVMe engine delivers superior write performance and more stable read throughput, surpassing the original SSD offloading technique.
@@ VI-B2 Assessment of the Integrated Solution
The end-to-end performance of MemAscend is evaluated by comparing its training throughput against the ZeRO-Infinity baseline under two hardware configurations described in Table II. Both systems were tested with the direct NVMe engine enabled. ZeRO-Infinity without this engine, which relies on a traditional filesystem-based approach, is unstable and prone to hanging, so it was excluded from this comparison. Table III reports the percentages of improvement in performance in a range of models, batch sizes, and hardware configurations, along with the gains achieved by MemAscend over ZeRO-Infinity. All tests used a fixed context length of 4096. For each configuration, we include several batch sizes to capture workload variation. The table lists the model, the batch sizes used for Configurations 1 and 2, and the corresponding percentage improvements.
TABLE III: End-to-End performance improvement from ZeRO-Infinity to MemAscend for Configurations 1 (C1) and 2 (C2)
@@ VI-B3 Extended Experiments
This section presents additional experiments that fall outside the main scope of the paper. The first explores the use of bf16 precision for optimizer-related values. This setup differs from the baseline, which relies on fp32 for optimizer-related values as part of fp16 mixed-precision training with SSD offloading. Because of this difference in numerical precision, the results are not directly comparable to those of ZeRO-Infinity. Nonetheless, enabling bf16 for optimizer states shows promising improvements, especially in reducing I/O load in SSD offloading systems. The second subsection examines mixed-precision training with bf16 instead of the fp16-based setup used in the main experiments. This test demonstrates the method’s adaptability and effectiveness in a bf16 training environment. Together, these experiments highlight MemAscend’s flexibility and stability across different precisions.
@@
The original baseline supported only fp32 optimizer states. Recent work has investigated lower-precision alternatives, including 8-bit [20] and 4-bit [21] formats. These reduced-precision representations are especially useful in SSD offloading systems, where lowering the I/O demand is critical. To explore this benefit, bf16 optimizer support is integrated into MemAscend. Then, the I/O volume per iteration (Figure 19) and end-to-end throughput improvements (Table IV) are evaluated on both Configurations 1 and 2, using a fixed context length of 4096 across four models. The evaluation results show that integrating bf16 into MemAscend consistently improves throughput across all tested models and configurations. For Configuration 1, the average throughput gain is 27.25%, peaking at 56.80% for Qwen2.5-7B with a batch size of 8, and dipping to 13.24% for Llama3.1-8B with a batch size of 80. Configuration 2 shows an average improvement of 17.08%, ranging from 9.99% for Qwen2.5-7B at batch size 20 to 24.21% for Qwen2.5-32B at batch size 4. These performance gains stem from lower I/O overhead, as shown in Figure 19, where bf16 usage significantly reduces I/O volume per iteration. The improvements are more pronounced at smaller batch sizes, where I/O bottlenecks dominate. In contrast, larger batches increase computation time, masking I/O savings. These results underscore the effectiveness of bf16 in optimizing SSD offloading systems.
@@
TABLE IV: Throughput improvement from using bf16 optimizer in MemAscend for Configurations 1 (C1) and 2 (C2)
@@ VII Conclusion
MemAscend re-frames SSD-offloading-based fine-tuning by showing that system memory, rather than GPU capacity, is the decisive constraint on commodity workstations. A detailed analysis uncovers four hidden overheads in the offloading pipeline: buffer-pool fragmentation, over-aligned pinned allocations, overflow-check spikes, and filesystem I/O overhead. MemAscend neutralizes them with the adaptive buffer pool, the alignment-free pinned memory allocation, the fused overflow check mechanism, and the direct NVMe engine for system memory reduction. Together, these techniques cut parameter buffer waste by up to 72.7%, reduce peak system memory demand by 55.7%, eliminate the 1.25× overflow check headroom, and improve throughput by up to 24.21% and 56.80% on different hardware configurations when bf16 half-precision optimizer is enabled. The recovered memory headroom immediately translates into support for larger models, longer context windows, and higher batch sizes on unchanged hardware, making full-parameter LLM fine-tuning genuinely accessible to smaller research groups and independent practitioners.

### v3

@@ V Memory Efficiency and Performance Analysis
By streamlining memory consumption, the MemAscend framework reduces system memory demand relative to earlier SSD offloading techniques. Those lower system memory footprints can translate directly into cheaper training setups; in other words, the same hardware budget can now accommodate models with more trainable parameters. The lower system memory usage also unlocks longer context windows, because the larger activation-checkpoint tensors required for extended sequences can be moved from GPUs to system memory without exceeding capacity. Therefore, instead of slowing down the training process, the overall performance benefits follow as well. Meanwhile, the extra system memory headroom permits larger batch sizes, driving GPUs toward higher utilization and boosting overall throughput. The upcoming analysis examines these gains in detail and explains how the reclaimed RAM supports longer context lengths and bigger batches, both of which contribute to faster training at lower cost. These experiments are conducted using Configuration 1, as described in Table III.
@@ V-A Memory Efficiency
Figure 10 first breaks down system memory usage of the proposed MemAscend while training the Qwen2.5-7B model [34], highlighting both the individual contributions and the synergistic effects of its four optimization components. Notably, both the compared Zero-Infinity and the proposed MemAscend are integrated with the Liger-Kernel, Flash-Attention, and offloaded gradient checkpointing for reducing residual memory usage. First, the adaptive buffer pool reduces the size of the parameter buffer pool by alleviating fragmentation issues and peak buffer usage (for example, from 9.14 GiB to 2.46 GiB, a 73% reduction), thereby lowering the baseline memory that must be pinned and streamed from the SSD. Second, the alignment-free pinned memory allocation reduces pinned memory overhead (from 24.90 GiB to 1.63 GiB), further decreasing memory usage. This not only resolves the overhead from large components like the gradient flat buffer but also improves the pinned memory overhead caused by the adaptive buffer pool, since it also uses pinned memory for allocation. Third, the direct NVMe engine eliminates filesystem serialization, improving performance for all SSD I/O, including the adaptive buffer pool buffers. Finally, the fused overflow check mechanism removes intermediate buffers, avoiding 35.46 GiB of overflow check overhead. Notably, since each component alleviates bottlenecks that would otherwise reduce the effectiveness of the others, the four optimizations together achieve significantly greater memory efficiency and performance improvements than any one component alone.
In Figure 10, it can be observed that the power-of-two alignment policy in conventional pinned memory allocation can nearly double memory usage, resulting in 24.90 GiB of allocator-induced overhead. By contrast, MemAscend’s alignment-free pinned memory allocation matches the requested tensor size precisely, lowering this overhead to just 1.63 GiB and achieving a 93.5% reduction in unnecessary memory consumption. Meanwhile, some components remain constant across methods, including the gradient flat buffer and small system allocations like optimizer state buffers and the swap-out buffer. These account for stable memory usage of roughly 28.37 GiB and 11.17 GiB, respectively.
On the other hand, to understand the theoretical minimal system memory that would still sustain equivalent throughput, Figure 10 presents the lower-bound memory requirement for ZeRO-Infinity and MemAscend. In this theoretical case, only the parameter buffer pool and the gradient flat buffer are strictly necessary: the former for streaming weights during training and the latter to avoid excessive I/O that would harm throughput. All framework-induced overheads—pinned memory, fragmentation, auxiliary allocations—are excluded. Based on this baseline, MemAscend lowers peak memory from 109.04 GiB (ZeRO-Infinity) to 43.64 GiB (60% reduction), significantly narrowing the gap to the theoretical minimum. The remaining margin for MemAscend is 12.81 GiB (29%), while ZeRO-Infinity still requires a 72% reduction to match this bound. This reduction improves scalability to larger models and longer sequence lengths under fixed memory budgets.
@@ VI-A Setup
The detailed hardware specifications for running the experiments are provided in Table III. MemAscend is evaluated using mixed-precision training of large language models with SSD offloading. The workloads vary by model architecture, context length, and batch size. The models include Llama3.1-8B and Qwen2.5, with parameter sizes of 7 billion, 14 billion, and 32 billion for most of our experiments. To test the diversity of model architectures, the model Qwen3-30B-A3B is also included to demonstrate a modern sparse language model. All models use the Hugging Face Transformers library (v4.51.0). The context lengths range from 4,096 to 131,072 tokens. Batch sizes vary across different experiments. The main evaluation metric is peak system memory usage during training and training throughput in tokens per second to assess the overall performance.
@@
Figure 13: Comparison of memory overhead for overflow checks between ZeRO-Infinity and MemAscend across various models.
TABLE III: Two Hardware Specifications
Component Configuration 1 Configuration 2
CPU Intel Xeon 6780E 2 $\times$ AMD EPYC 7282
CPU Memory 1 TB 6400 MT/s DDR5 1 TB 3200 MT/s DDR4
PCIe Gen5 Gen4
GPU 2 $\times$ NVIDIA H100 PCIe (w/o NVL) 1 $\times$ NVIDIA A5000
SSD 1 $\times$ DapuStor Haishen5 H5100 (7.5 TB) 2 $\times$ Phison AI100E
To evaluate the effectiveness of MemAscend, the SOTA ZeRO-Infinity is compared, and both systems share a common setup to ensure fairness. Static memory components, including model parameters and optimizer states, are stored entirely on the SSD. To reduce GPU memory usage, gradient checkpointing, Flash-Attention, and the Liger-Kernel are enabled for both setups. For experiments that require offloaded gradient checkpointing, a modified version of Unsloth is utilized, adapted from its original single-GPU design in which the original Unsloth implementation of offloaded activation checkpointing only supported a single-GPU setup and hardcoded the CUDA device index in its source code. The modification replaces the hardcoded index with the correct device index for each process in a multi-GPU scenario.
VI-B Experimental Results
@@ VI-B1 Assessment of Individual Methods
This evaluation examines the performance of the direct NVMe engine, described in Section IV-E, for SSD I/O in training scenarios with SSD offloading. The engine manages offloading of fp16 and fp32 model weights, gradients, and optimizer states. To determine suitable tensor sizes for benchmarking, model weight distributions were analyzed, excluding tensors that are too small to benefit from offloading. For instance, tensors with fewer than two million elements perform better in CPU memory. Common dense LLM architectures include several layers suitable for SSD offloading, such as feedforward layers, q, k, v, and o projection layers, the embedding layer, and the linear head layer. To ensure a precise comparison with the original SSD offloading technique, tensor size ranges for fp16 and fp32 data were collected across various models. The analysis also considers multi-GPU setups, where data are split across processes. For example, a 100-million-element tensor divides into 50 million elements per process in a two-GPU configuration. Benchmarks used Configuration 2, as specified in Table III. The original SSD offloading system, reliant on filesystem I/O, utilized two SSDs in a RAID0 array formatted with ext4. By contrast, MemAscend employs the eirect NVMe engine, which bypasses software RAID and sends requests directly to SSDs through the NVMe driver.
Figures 14(a) and 14(b) demonstrate that the direct NVMe engine significantly lowers write latency and improves write bandwidth across different transfer sizes. For a 2,097,152-byte tensor, latency decreases from 988.222 microseconds to 219.418 microseconds. For a 3,114,270,720-byte tensor, it decreases from 304,604.779 microseconds to 266,212.637 microseconds. The bandwidth improvement reaches up to 4.5×, with a 72.04% average gain. The decreasing trend in MemAscend’s write bandwidth stems from the SSD’s internal DRAM buffers and SLC write cache, which can absorb short bursts. When the transfer size is small and the I/O queue is efficiently utilized, data is served directly from high-speed cache, yielding very high application-observed bandwidth. As the written volume grows, these caches fill, forcing data to be destaged to slower NAND media. The resulting throughput gradually converges toward the sustained rate of the device. In contrast, the filesystem-based baseline performs poorly for small writes due to dominant host-side overheads, then improves as these costs are amortized, leading to the contrasting shapes of the two curves.
The read behavior differs, as shown in Figures 14(c) and 14(d). Reads bypass much of the expensive journaling and allocation paths, so both methods achieve similar average read bandwidth. However, MemAscend’s direct LBA I/O avoids pathname resolution, metadata lookups, and RAID-level merges on the critical path, thereby reducing variability. This results in more consistent read performance, even when the mean bandwidth remains comparable to the baseline.
@@ VI-B2 Assessment of the Integrated Solution
TABLE IV: End-to-End performance improvement from ZeRO-Infinity to MemAscend for Configurations 1 (C1) and 2 (C2)
@@
The end-to-end performance of MemAscend is evaluated by comparing its training throughput against the ZeRO-Infinity baseline under two hardware configurations described in Table III. Both systems were tested with the direct NVMe engine enabled. ZeRO-Infinity without this engine, which relies on a traditional filesystem-based approach, is unstable and prone to hanging, so it was excluded from this comparison. Table IV reports the percentages of improvement in performance in a range of models, batch sizes, and hardware configurations, along with the gains achieved by MemAscend over ZeRO-Infinity. All tests used a fixed context length of 4096. For each configuration, we include several batch sizes to capture workload variation. The table lists the model, the batch sizes used for Configurations 1 and 2, and the corresponding percentage improvements.
Performance Across Diverse Model Architectures
TABLE V: Hardware Specifications for Sparse Model (MoE)
Component Configuration 3
CPU Intel Xeon 8480+
CPU Memory 1 TB 4800 MT/s DDR5
PCIe Gen5
GPU 2 $\times$ NVIDIA H100 SXM5 (with NVL)
SSD 2 $\times$ Samsung 980 Pro (1 TB each)
To demonstrate the memory efficiency of MemAscend on diverse model architectures, this section evaluates its performance on sparse models, also known as the Mixture-of-Experts (MoE) model (Configuration 3 in Table V). Figures 18(a) and 18(b) show that baseline memory usage grows from 756.73 GiB at 4,096 tokens to 818.74 GiB at 131,072 tokens, whereas MemAscend increases only from 202.24 GiB to 248.75 GiB, achieving average reductions of 71.87% and 71.40%, respectively. MoE architectures introduce many small expert feed-forward modules (up to 128 for Qwen3-30B-A3B), increasing the number of buffers needed for SSD offloading. MemAscend’s adaptive buffer pool mitigates this pressure by allocating only the exact buffer size required for each expert. In contrast, the baseline allocator must size all buffers to match the largest layer (e.g., embeddings), resulting in significant unnecessary overhead. These results confirm that MemAscend maintains strong memory efficiency and robustness across both dense and sparse model families.
(a) Different Context Length
(b) Different Batch Size
Figure 18: Comparison of system memory usage and throughput between ZeRO-Infinity and MemAscend across different context lengths (batch size = 1) and batch sizes (context length = 4096).
Convergence Behavior
To understand whether the proposed MemAscend changes convergence behavior, we compare its loss trajectory against ZeRO-Infinity using a real training workload. MemAscend introduces only system-level optimizations—buffer management, pinned-memory allocation, fused overflow checking, and direct NVMe access—without modifying numerical kernels. Therefore, no accuracy deviation is expected. This study evaluate Qwen2.5-0.5B on the OpenWebText dataset [35] with batch size 128 and sequence length 1024 for 250 training steps. Figures 19(a) and 19(b) plot the training and evaluation loss curves, respectively. As shown, MemAscend exhibits identical convergence behavior to ZeRO-Infinity throughout the entire training trajectory, confirming that our framework fully preserves model accuracy.
(a) Training
(b) Validation
Figure 19: Comparison of loss convergence behavior between ZeRO-Infinity and MemAscend.
@@ VI-B3 Extended Experiments
This section presents additional experiments that fall outside the main scope of the paper. The first explores the use of bf16 precision for optimizer-related values. This setup differs from the baseline, which relies on fp32 for optimizer-related values as part of fp16 mixed-precision training with SSD offloading. Because of the difference in numerical precision, the loss convergence behavior will be different compared to ZeRO-Infinity. Nonetheless, enabling bf16 for optimizer states shows promising improvements, especially in reducing I/O load in SSD offloading systems. The second subsection examines mixed-precision training with bf16 instead of the fp16-based setup used in the main experiments. This test demonstrates the method’s adaptability and effectiveness in a bf16 training environment. Together, these experiments highlight MemAscend’s flexibility and stability across different precisions.
@@
The original baseline supported only fp32 optimizer states. Recent work has investigated lower-precision alternatives, including 8-bit [20] and 4-bit [21] formats. These reduced-precision representations are especially useful in SSD offloading systems, where lowering the I/O demand is critical. To explore this benefit, bf16 optimizer support is integrated into MemAscend. Notably, bf16 is selected for optimizer state compression because it incurs minimal numerical-precision overhead and requires only direct truncation from fp32. More aggressive formats (e.g., fp8, 8-bit or 4-bit quantization) require additional scaling, dynamic-range control, and overflow handling, which would introduce added computational complexity not central to the focus of this study. Then, the I/O volume per iteration (Figure 21) and end-to-end throughput improvements (Table VI) are evaluated on both Configurations 1 and 2, using a fixed context length of 4096 across four models. The evaluation results show that integrating bf16 into MemAscend consistently improves throughput across all tested models and configurations. For Configuration 1, the average throughput gain is 27.25%, peaking at 56.80% for Qwen2.5-7B with a batch size of 8, and dipping to 13.24% for Llama3.1-8B with a batch size of 80. Configuration 2 shows an average improvement of 17.08%, ranging from 9.99% for Qwen2.5-7B at batch size 20 to 24.21% for Qwen2.5-32B at batch size 4. These performance gains stem from lower I/O overhead, as shown in Figure 21, where bf16 usage significantly reduces I/O volume per iteration. The improvements are more pronounced at smaller batch sizes, where I/O bottlenecks dominate. In contrast, larger batches increase computation time, masking I/O savings. These results underscore the effectiveness of bf16 in optimizing SSD offloading systems.
@@
TABLE VI: Throughput improvement from using bf16 optimizer in MemAscend for Configurations 1 (C1) and 2 (C2)
@@ VII Conclusion
MemAscend re-frames SSD-offloading-based fine-tuning by showing that system memory, rather than GPU capacity, is the decisive constraint on commodity workstations. A detailed analysis uncovers four hidden overheads in the offloading pipeline: buffer-pool fragmentation, over-aligned pinned allocations, overflow-check spikes, and filesystem I/O overhead. MemAscend neutralizes them with the adaptive buffer pool, the alignment-free pinned memory allocation, the fused overflow check mechanism, and the direct NVMe engine for system memory reduction. Together, these techniques cut parameter buffer waste by up to 72.7%, reduce peak system memory demand by 55.7%, eliminate the 1.25× overflow check headroom, if users can tolerate a slight loss of precision in optimizer states, MemAscend can improve throughput by up to 56.80% and 24.21% on two different hardware configurations when the bf16 half-precision optimizer is enabled. The recovered memory headroom immediately translates into support for larger models, longer context windows, and higher batch sizes on unchanged hardware, making full-parameter LLM fine-tuning genuinely accessible to smaller research groups and independent practitioners.